# Prototype-Guided Diffusion: Visual Conditioning without External Memory

### v3

@@ IV Method
Input: Dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, number of prototypes $K$, noise schedule $\{\alpha_{t}\}_{t=1}^{T}$, weights $\tau,\alpha,\beta$
@@ 24  Combine losses:
25 $\mathcal{L}_{\text{PDM}}=\mathcal{L}_{\text{diff}}+\mathcal{L}_{\text{contrastive}}+\alpha\cdot\mathcal{L}_{\text{align}}+\beta\cdot\mathcal{L}_{\text{compact}}$;
@@ V-B Ablation Study: Number of Prototypes in PDM
Table III shows that reducing the number of prototypes below the number of classes leads to a clear drop in performance across all metrics. For example, decreasing from $10$ to $3$ prototypes reduces IS from $8.35$ to $7.50$ and increases FID from $8.10$ to $11.90$, indicating lower generation quality and poorer alignment with the real data distribution. This degradation occurs because fewer prototypes cannot represent the semantic diversity of the dataset, forcing multiple classes to share a prototype and thereby merging distinct modes in the data distribution.
When increasing the number of prototypes beyond $10$, IS improves slightly up to $15$ prototypes, reflecting a modest gain in sample diversity. However, FID and KID also decrease slightly, suggesting better fidelity and distributional alignment. At $20$ prototypes, IS drops again, likely due to over-fragmentation of the latent space: too many prototypes may split coherent clusters, leading to redundant or less meaningful centroids without adding new semantic structure.

### v4

@@ IV Method
Input: Dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, number of prototypes $K$, noise schedule $\{\alpha_{t}\}_{t=1}^{T}$, weight $\tau$
@@ 24  Combine losses:
25 $\mathcal{L}_{\text{PDM}}=\mathcal{L}_{\text{diff}}+\mathcal{L}_{\text{contrastive}}+\mathcal{L}_{\text{align}}+\mathcal{L}_{\text{compact}}$;
@@ V-B Ablation Study: Number of Prototypes in PDM
Table III shows that reducing the number of prototypes below the number of classes leads to a clear drop in performance across all metrics. For example, decreasing from $10$ to 33 prototypes reduces IS from $8.35$ to $7.50$ and increases FID from $8.10$ to $11.90$, indicating lower generation quality and poorer alignment with the real data distribution. This degradation occurs because fewer prototypes cannot represent the semantic diversity of the dataset, forcing multiple classes to share a prototype and thereby merging distinct modes in the data distribution.
When increasing the number of prototypes beyond $10$, IS improves slightly up to $15$ prototypes, reflecting a modest gain in sample diversity. However, FID and KID also decrease slightly, suggesting better fidelity and distributional alignment. At 2020 prototypes, IS drops again, likely due to over-fragmentation of the latent space: too many prototypes may split coherent clusters, leading to redundant or less meaningful centroids without adding new semantic structure.